OpenStudy (anonymous):

a box contains five one dollar bills, nine five dollar bills, ten ten dollar bills, and nine twenty dollar bills. what is the expectation if one bill is selected. $10.00 $9.00 $11.00 $12.00

OpenStudy (michele_laino):

here we have to compute this: \[{\text{expectation}} = \frac{{5 + 5 \times 9 + 10 \times 10 + 9 \times 20}}{{33}} = \frac{{330}}{{33}} = 10\]